# Textual Entailment and Token Probability as Bias Evaluation Metrics

### v1

@@ Abstract
Measurement of social bias in language models is typically by token probability (TP) metrics, which are broadly applicable but have been criticized for their distance from real-world langugage model use cases and harms. In this work, we test natural language inference (NLI) as a more realistic alternative bias metric. We show that, curiously, NLI and TP bias evaluation behave substantially differently, with very low correlation among different NLI metrics and between NLI and TP metrics. We find that NLI metrics are more likely to detect “underdebiased” cases. However, NLI metrics seem to be more brittle and sensitive to wording of counterstereotypical sentences than TP approaches. We conclude that neither token probability nor natural language inference is a “better” bias metric in all cases, and we recommend a combination of TP, NLI, and downstream bias evaluations to ensure comprehensive evaluation of language models.
@@
Textual Entailment and Token Probability as Bias Evaluation Metrics
@@ 1 Introduction
For this reason, fairness experts recommend in situ evaluation of LM systems on realistic inputs, focusing on the salient risks of social bias for a system’s user base, domain, and intended purpose, such as the localized bias evaluation proposed by Pang et al. (2025). However, downstream bias evaluation is ill-suited for comparing the risk of social biases across a variety of LMs. Because in situ evaluation usually occurs on a model that has already been finetuned for a specific task, it is generally impractical to finetune multiple models to determine their relative safety risks. LM system designers who are choosing between possible base LMs and want to choose a base model that will minimize biases relevant to their specific users need generalizable, multiple-model bias evaluation.
Thus, there is a necessity for midstream bias evaluation metrics that evaluate an LM’s degree of social bias on realistic NLP tasks, but can be relatively easily applied to multiple models and do not rely on knowledge of a specific use case. We propose using natural language inference (NLI) — specifically, textual entailment — as a midstream bias evaluation task, which is both more realistic than TP metrics and more generalizable than customized in situ risk evaluation. The primary contributions of this work are:
  • Development and release of a novel NLI bias benchmark dataset [^1] Dataset and code available at https://github.com/katyfelkner/wq-nli.
@@
Through detailed analysis of the behavior of NLI and TP bias evaluations at multiple levels (stereotype categories, specific stereotypes, and individual test instances) and across three model families, nine models, and three debiasing conditions for each, we find significant differences in bias evaluation results. Crucially, we compare the two tasks on exactly the same set of community-sourced bias definitions, so any difference in evaluation results is due to the design of bias metrics, not the content of benchmark datasets.
@@ 2.1 Dataset Construction
In this work, we first create an NLI version of an existing bias benchmark dataset, which yields, to our knowledge, the first pair of bias datasets using different tasks but exactly the same bias definitions. We chose to use the WinoQueer (WQ) dataset Felkner et al. (2023) because of of its thoroughness and grounding. WQ is available under an MIT License, and our use is consistent with its intended use. WQ consists of 46,036 sentence pairs covering nine LGBTQ+ subgroups, four counterfactual groups, 173 unique attested harm predicates, and 19 template sentences. The predicates were sourced from a large-scale survey of the LGBTQ+ community and manually annotated by community members. In the original WQ dataset, bias scores are calculated from token probabilities, following the methodology established in Nangia et al. (2020). This method defines the bias score as the percentage of cases where the stereotypical model output has higher aggregated token probability than the counterstereotypical model output. These percentile bias scores are easy to interpret: 50% is the ideal score, and scores over 50 indicate the presence of stereotypical bias.
We introduce WinoQueer-NLI (WQ-NLI), a version of the WQ dataset that evaluates bias on a textual entailment classification task, instead of token probabilities. In this dataset, we preserve the broad coverage and community-sourced nature of the WQ dataset, as well as the intuitive interpretability of percentile bias scores; we also address the grounding and realism weaknesses of token probability-based bias evaluation by evaluating on textual entailment, a more realistic classification task. Textual entailment is a standard NLP task (e.g. Williams et al., 2018) in which the goal is to predict whether a hypothesis is entailed ($E$), contradicted ($C$), or neutral ($N$), given that a premise is true. NLI datasets consist of premise/hypothesis sentence pairs, and classifiers usually output a distribution $[p(E),p(N),p(C)]$.
In the bias evaluation setting, we consider the probability that a premise specifying an identity entails a hypothesis containing a harmful stereotype. The unbiased answer in all cases is neutral, following Dev et al. (2020), since the subject’s identity should provide no information on whether or not the stereotype is true; any conclusion otherwise is an indication of bias. A classification of ‘contradict’ signifies a tendency to apply attested stereotypes to non-targeted majority group. While this is less desirable than ’neutral,’ it is still considered preferable to ’entail,’ which would signify a model’s reinforcement of established harmful stereotypes. While the original token probability WinoQueer dataset (which we will henceforth call WQ-TP for clarity) consists of sentence pairs, WQ-NLI consists of sentence triples: a stereotypical premise sentence with a minority identity, a counterstereotypical premise sentence with a majority (counterfactual) identity, and a shared hypothesis sentence containing the actual stereotype.
@@
The similarity of the WQ-TP and WQ-NLI datasets enables meaningful direct comparison between token probability and textual entailment as LM bias metrics. WQ-NLI is the first community-grounded bias dataset using the textual entailment task. Most importantly, to the best of our knowledge, WQ-TP and WQ-NLI form the first directly comparable pair of bias datasets using the exact same set of bias definitions on two different evaluation tasks. WQ-NLI is released under an MIT license.
@@ 2.2 Model List and Compute Estimate
We experiment on nine models, listed with their number of parameters: BERT base cased (109M), BERT base uncased (110M), BERT large cased (335M), BERT large uncased (336M) (Devlin et al., 2019), RoBERTa base (125M), RoBERTa large (561M) (Liu et al., 2019), GPT2 (137M), GPT2 Medium (380M), and GPT2 XL (1.61B) (Radford et al., 2019). These are a subset of the models evaluated in Felkner et al. (2023), in which the authors trained two debiased versions of each tested model by continued pretraining on large corpora of community data. We chose these models in order to minimize our compute usage by starting from already-debiased models and finetuning for the NLI task. Therefore, for each of our nine models, we include three variants: raw, with no debiasing; news, which was debiased on mainstream news data about the relevant community; and twitter, which was debiased on social media data directly from the relevant community. Across experimentation, task finetuning, and evaluation, we used around 1,200 GPU-hours across NVIDIA P100, V100, and A40 GPUs.
@@ 2.3 MNLI Task Finetuning
Before evaluation on NLI bias metrics, all models are finetuned for the textual entailment task on the Multi-Genre Natural Language Inference (MNLI) dataset (Williams et al., 2018), a crowd-sourced textual entailment dataset containing about 400,000 examples. Each instance contains: (1) Premise: A sentence taken from existing text sources; (2) Hypothesis: A sentence written by human annotators to pair with the premise; (3) Label: The relationship between the premise and hypothesis—entailment, contradiction, or neutral. Following standard procedures, we train a linear classifier layer on top of each model and finetune all parameters of the model. For debiased models, task finetuning is done after debiasing. All models are finetuned for four epochs on the MNLI training set, and all reach accuracy scores comparable with published MNLI results for the same models. We conduct one finetuning run for each model.
@@ 2.4 NLI as an Evaluation Metric
To use trained NLI models as a bias evaluation task, we collect two probability triples, one for the stereotypical premise and one for the counterstereotypical premise, each using the same shared hypothesis. Thus, for every triple of sentences in the WQ-NLI benchmark dataset, our evaluation results in a sextuple of probabilities: [p(E|S),p(N|S),p(C|S),p(E|S~),p(N|S~),p(C|S~)][$p(E|\tilde{S}),p(N|\tilde{S}),p(C|\tilde{S})$], where $S$ is the stereotypical premise and $\tilde{S}$ is the counterstereotypical premise.
One of the desirable properties of token probability bias evaluation metrics is the intuitiveness of their percentile bias scores, which are calculated as the percentage of test instances on which the model displays stereotypical bias, as defined by a comparison of summed pseduo-log-probabilities for the stereotypical and counterstereotypical sentences in each pair. Thus, we explore several options for pairwise comparison that would transform the set of six raw probabilities from NLI evaluation into a similar aggregate percentile bias score.
@@
$M_{7}$ $\operatorname*{argmax}(p(x|S))=E\wedge\operatorname*{argmax}(p(x|\tilde{S}))=N$
$M_{8}$ $\operatorname*{argmax}(p(x|S))=E\wedge\operatorname*{argmax}(p(x|\tilde{S}))=C$
@@
However, these approaches ignore the difference between neutrality and contradiction, which is conceptually significant when quantifying a model’s implicit associations between demographic groups and harmful stereotypes. This additional granularity of output is one of the appeals of NLI over token probability as a bias evaluation task. Therefore, we test several other metrics that take into account neutral and contradict probabilities as well as the entailment probability. $M_{5}$ is designed to penalize a model for assigning probability to the attested harm being entailed, while giving “credit” for assigning probability to contradiction. $M_{6}$ is similar to $M_{5}$, but it gives “full credit” for contradict probability and “half credit” for neutral probability. $M_{7}$ and $M_{8}$ consider the model’s highest probability outcome for both stereotypical and counterstereotypical test pairs. $M_{7}$ counts the instances where the stereotypical pair is most likely entailed and the counterstereotypical pair is most likely neutral, i.e. cases where the model associates the minority identity with the stereotype being true but does not associate the majority identity with the stereotype being either true or false. Similarly, $M_{8}$ counts instances where the stereotypical pair is most likely entailed and the counterstereotypical pair is most likely contradicted, i.e. cases where the model associates the minority identity with the stereotype being true and associates the majority identity with the stereotype being false.
To select a conversion from entailment probabilities to percentile bias scores, we first examine the $R^{2}$ values for the correlation between token probability bias scores and NLI bias scores for each tested model. The results of this analysis are described in Section 3.2. In general, we find at best weak correlation between TP and NLI metrics or among the tested NLI metrics. To better understand the reasons for this behavior, we conduct several fine-grained analyses of the behavior of TP and NLI as bias evaluation tasks. To facilitate this analysis, we manually code the attested harm predicates from the original WQ dataset into 18 categories, which are listed in Appendix A.1.
@@ 3.1 WQ-NLI Baseline Results
Table 3 shows the WQ-TP and WQ-NLI bias scores for raw and debiased models. The key takeaway from these results is that TP bias scores are not a reliable predictor of NLI bias scores. Therefore, a model which scores well on a token probability bias evaluation may still display social biases in a downstream bias evaluation (such as WQ-NLI) or in a deployment context. Additionally, there are several cases (BERT Large, RoBERTa, GPT2) where models appear to be debiased according to the TP evaluation, but actually perform similarly or worse on the NLI evaluation. This suggests that token probability alone is not sufficient to determine whether a model has been sufficiently debiased.
Model TP Raw TP News TP Twitter NLI Raw NLI News NLI Twitter
BERT Base Uncased 74.49 45.71 41.05 62.53 51.27 51.43
BERT Base Cased 64.40 61.67 57.81 51.26 53.33 49.28
BERT Large Uncased 64.14 53.10 43.19 69.41 63.92 65.48
BERT Large Cased 70.69 58.52 56.94 63.22 79.50 70.68
RoBERTa Base 69.18 64.33 54.34 73.48 77.21 77.65
RoBERTa Large 71.09 57.19 58.45 72.64 72.94 65.36
GPT2 68.27 49.82 45.11 57.91 52.91 56.18
GPT2 Medium 55.83 44.29 38.73 59.93 60.52 72.39
GPT2 XL 66.15 65.33 36.73 60.50 58.06 48.31

### v2

# Textual Entailment is not a Better Bias Metric than Token Probability

@@ Abstract
Measurement of social bias in language models is typically by token probability (TP) metrics, which are broadly applicable but have been criticized for their distance from real-world language model use cases and harms. In this work, we test natural language inference (NLI) as an alternative bias metric. In extensive experiments across seven LM families, we show that NLI and TP bias evaluation behave substantially differently, with very low correlation among different NLI metrics and between NLI and TP metrics. NLI metrics are more brittle and unstable, slightly less sensitive to wording of counterstereotypical sentences, and slightly more sensitive to wording of tested stereotypes than TP approaches. Given this conflicting evidence, we conclude that neither token probability nor natural language inference is a “better” bias metric in all cases. We do not find sufficient evidence to justify NLI as a complete replacement for TP metrics in bias evaluation.
@@
Textual Entailment is not a Better Bias Metric than Token Probability
@@ 1 Introduction
For this reason, fairness experts recommend in situ evaluation of LM systems on realistic inputs, such as the localized bias evaluation proposed by Pang et al. (2025). However, downstream bias evaluation is ill-suited for comparing the risk of social biases across a variety of LMs. Because such evaluation usually occurs on a model that has already been finetuned for a specific task, it is generally impractical to finetune multiple models to determine their relative safety risks. LM system designers who are choosing between possible base LMs and want to choose a base model that will minimize biases relevant to their specific users need generalizable, multiple-model bias evaluation.
Thus, there is a necessity for alternatives to TP bias evaluation metrics, which should be easily applicable to multiple models and should not rely on knowledge of a specific use case. Ideally, this metric should also be somewhat reflective of real NLP tasks, while maintaining general applicability. We propose using natural language inference (NLI) — specifically, textual entailment classification — as an alternative bias evaluation task. The primary contributions of this work are:
  • Development and release of a novel NLI bias benchmark dataset [^1] Dataset and code available at https://github.com/katyfelkner/wq-nli and are released under MIT License.
@@
Through detailed analysis of the behavior of NLI and TP bias evaluations at multiple levels (stereotype categories, specific stereotypes, and individual test instances) and across seven model families, seventeen models, and three debiasing conditions, we find significant differences in bias evaluation results. Crucially, we compare the two tasks on exactly the same set of community-sourced bias definitions, so any difference in evaluation results is due to the design of bias metrics, not the content of benchmark datasets. We find that TP and NLI behave differently as bias metrics, but there is insuffienct evidence that NLI is suitable as a drop-in replacement for TP bias metrics.
@@ 2.1 Dataset Construction
In this work, we first create an NLI version of an existing bias benchmark dataset, which yields, to our knowledge, the first pair of bias datasets using different tasks but exactly the same bias definitions. We chose to use the WinoQueer (WQ) dataset Felkner et al. (2023) because of its thoroughness and grounding. [^2] WQ is available under an MIT License, and our use is consistent with its intended use. WQ consists of 46,036 sentence pairs covering nine LGBTQ+ subgroups, four counterfactual groups, 173 unique attested harm predicates, and 19 template sentences. The predicates were sourced from a large-scale survey of the LGBTQ+ community and manually annotated by community members. In the original WQ dataset (which we will henceforth call WQ-TP for clarity), bias scores are calculated from token probabilities, following the methodology established in Nangia et al. (2020). This method defines the bias score as the percentage of cases where the stereotypical model output has higher aggregated token probability than the counterstereotypical model output, producing interpretable percentile bias scores where scores over 50% indicate a biased model.
We introduce WinoQueer-NLI (WQ-NLI), a version of the WQ dataset that evaluates bias on a textual entailment classification task, instead of token probabilities. Textual entailment is a standard NLP task (e.g. Williams et al., 2018) in which the goal is to predict whether a hypothesis is entailed ($E$), contradicted ($C$), or neutral ($N$), given that a premise is true. NLI datasets consist of premise/hypothesis sentence pairs, and classifiers usually output a distribution $[p(E),p(N),p(C)]$.
In the bias evaluation setting, we consider the probability that a premise specifying an identity entails a hypothesis containing a harmful stereotype. The unbiased answer in all cases is neutral, following Dev et al. (2020), since the subject’s identity should provide no information on whether or not the stereotype is true. Any conclusion otherwise is an indication of bias. A classification of ‘contradict’ signifies a tendency to apply attested stereotypes to non-targeted majority group. While this is less desirable than ‘neutral,’ it is still considered preferable to ‘entail,’ which would signify a model’s reinforcement of established harmful stereotypes. WQ-NLI consists of sentence triples: a stereotypical premise sentence with a minority identity, a counterstereotypical premise sentence with a majority (counterfactual) identity, and a shared hypothesis sentence containing the actual stereotype.
@@
To the best of our knowledge, WQ-TP and WQ-NLI form the first directly comparable pair of bias datasets using the exact same set of bias definitions on two different evaluation tasks. This similarity in bias benchmark content enables meaningful direct comparison between token probability and textual entailment as LM bias metrics.
@@ 2.2 Model List and Compute Estimate
We experiment on seventeen models across the BERT, RoBERTa, GPT2, LLAMA, Qwen, Gemma, and Mistral families. The BERT, RoBERTa, and GPT2 models included are a subset of the models evaluated in Felkner et al. (2023), in which the authors trained two debiased versions of each tested model by continued pretraining on large corpora of community data. We chose these models in order to minimize our compute usage by starting from already-debiased models and finetuning for the NLI task. Therefore, for each of these nine models, we include three variants: raw, with no debiasing; news, which was debiased on mainstream news data about the relevant community; and twitter, which was debiased on social media data directly from the relevant community. Because BERT, RoBERTa, and GPT2 are relatively old and relatively small, we also extend our model selection to include 8 modern LMs up to 8B parameters across the Llama 3, Qwen 3, Gemma, and Mistral families. However, due to compute requirements for continued pretraining of large models, we do not replicate the debiasing via community exposure procedure from Felkner et al. (2023) on these 8 newer models. Across experimentation, task finetuning, and evaluation, we used around 1,600 GPU-hours across NVIDIA P100, V100, and A40 GPUs.
@@ 2.3 MNLI Task Finetuning
Before evaluation on NLI bias metrics, all models are finetuned for the textual entailment task on the Multi-Genre Natural Language Inference (MNLI) dataset (Williams et al., 2018), a crowd-sourced textual entailment dataset containing about 400,000 examples. Following standard procedures, we train a linear classifier layer on top of each model and finetune the base model. For BERT, RoBERTa, and GPT2, we finetune all parameters of the base model. For Llama, Qwen, Gemma, and Mistral, we use Low-Rank Adaptation (LoRA) for parameter-efficient finetuning. For debiased models, task finetuning is done after debiasing. All models are finetuned for four epochs on the MNLI training set, and all reach accuracy scores comparable with published MNLI results for the same models. We conduct one finetuning run for each model.
@@ 2.4 NLI as an Evaluation Metric
For every triple of sentences in the WQ-NLI benchmark dataset, our evaluation results in a sextuple of probabilities: [p(E|S),p(N|S),p(C|S),p(E|S~),p(N|S~),p(C|S~)][$p(E|\tilde{S}),p(N|\tilde{S}),p(C|\tilde{S})$], where $S$ is the stereotypical premise and $\tilde{S}$ is the counterstereotypical premise.
One of the desirable properties of token probability bias evaluation metrics is the intuitiveness of their percentile bias scores, which are calculated as the percentage of test instances on which the model displays stereotypical bias, as defined by a comparison of summed pseudo-log-probabilities for the stereotypical and counterstereotypical sentences in each pair. Thus, we explore several options for pairwise comparison that would transform the set of six raw probabilities from NLI evaluation into a similar aggregate percentile bias score.
@@
[…]
$M_{10}$ $\operatorname*{argmax}(p(x|S))=E\wedge\operatorname*{argmax}(p(x|\tilde{S}))=N$
$M_{11}$ $\operatorname*{argmax}(p(x|S))=E\wedge\operatorname*{argmax}(p(x|\tilde{S}))=C$
@@
However, these approaches ignore NLI’s distinction between neutrality and contradiction. Therefore, we test several other metrics that take into account neutral and contradict probabilities as well as the entailment probability. $M_{5}$ is designed to penalize a model for assigning probability to the attested harm being entailed, while giving “credit” for assigning probability to contradiction. $M_{6}$ is similarly penalizes the model for entailment but gives “credit” for assigning probability to neutral. This closely mirrors the NLI task formulation, where the correct answer is always neutral and entailment is considered more harmful than contradiction. $M_{7}$ is similarly penalizes the model for entailment and gives equal “credit” for both neutral and contradiction. $M_{8}$ is similar to $M_{7}$, but it gives “full credit” for contradict probability and “half credit” for neutral probability. The intuition here is that actively contradicting a stereotype should be more heavily rewarded than a neutral conclusion. $M_{9}$ is similar to $M_{7}$ and is the opposite of $M_{8}$. It gives “full credit” for neutral probability and “half credit” for neutral probability. The intuition for $M_{9}$ is that neutral is the correct answer and should be heavily rewarded, but contradiction is preferable to entailment and should receive some reward.
$M_{10}$ and $M_{11}$ consider the model’s highest probability outcome for both stereotypical and counterstereotypical test pairs. $M_{10}$ counts the instances where the stereotypical pair is most likely entailed and the counterstereotypical pair is most likely neutral. Similarly, $M_{11}$ counts instances where the stereotypical pair is most likely entailed and the counterstereotypical pair is most likely contradicted.
To select a conversion from entailment probabilities to percentile bias scores, we first examine the $R^{2}$ values for the correlation between token probability bias scores and NLI bias scores for each tested model. The results of this analysis are described in Section 3.2. In general, we find at best weak correlation between TP and NLI metrics or among the tested NLI metrics. To better understand the reasons for this behavior, we conduct a mutual information analysis of the behavior of TP and NLI as bias evaluation tasks. To facilitate this analysis, we manually code the attested harm predicates from the original WQ dataset into 18 categories, which are listed in Appendix A.1.
@@ 3.1 WQ-NLI Baseline Results
Table 3 shows the WQ-TP and WQ-NLI bias scores for raw and debiased models. First, we observe that on WQ-TP, newer models are not necessarily less biased than older models, despite increased safety and alignment efforts in recent models. Llama 3, Gemma, and Mistral have bias scores roughly comparable to BERT and RoBERTa. Qwen 3 is a notable exception to this trend: on WQ-TP, the 1.7B and 8B models are very close to unbiased, and the 4B model has a lower bias score than most other models.
@@
The key takeaway from the WQ-NLI results is that TP bias scores are not a reliable predictor of NLI bias scores. Therefore, a model which scores well on a token probability bias evaluation may still display social biases in a task-based bias evaluation (such as WQ-NLI) or in a deployment context. Examples of this phenomenon (bold in Table 3) include GPT2 Medium, Qwen 3 1.7B, and Qwen 3 8B. Conversely, some models appear to be less biased according to NLI metrics but still show much more bias on TP bias metrics; examples (italicized in Table 3) include most BERT and GPT2 models, as well as Llama 3.1 8B and Gemma 7B.
Model TP $M_{1}$ $M_{2}$ $M_{3}$ $M_{4}$ $M_{5}$ $M_{6}$ $M_{7}$ $M_{8}$ $M_{9}$ $M_{10}$ $M_{11}$
BERT B. Un. 74.5 62.5 25.4 16.0 9.0 61.1 54.7 62.5 61.3 57.1 2.5 11.9
BERT B. C. 64.4 51.3 17.1 10.1 5.1 56.5 40.7 51.3 55.3 42.9 1.4 5.7
BERT Lg. Un. 64.1 69.4 11.3 7.3 5.6 67.7 43.5 69.4 67.7 45.9 2.2 4.7
BERT Lg. C. 70.7 63.2 12.1 6.9 2.4 62.3 46.4 63.2 62.4 48.7 2.4 4.2
RoBERTa B. 69.2 73.5 24.0 11.1 3.9 70.2 52.3 73.5 71.5 59.1 2.1 10.4
RoBERTa Lg. 71.1 72.6 14.8 11.3 8.7 76.5 36.9 72.6 76.6 41.4 2.2 8.6
GPT2 68.3 57.9 22.2 8.6 2.9 59.6 47.8 57.9 59.3 53.0 2.3 5.8
GPT2 Med. 55.8 59.9 11.2 5.2 3.1 67.0 39.7 59.9 66.4 45.3 0.7 4.1
GPT2 XL 66.2 60.5 6.9 6.9 6.9 67.5 50.7 66.5 68.7 56.3 1.2 5.1
Llama 3.2 1B 66.0 70.9 16.6 4.0 1.1 54.4 61.7 70.7 57.0 70.9 0.8 4.1
Llama 3.2 3B 67.0 67.3 10.7 5.7 3.9 67.6 44.3 67.9 68.6 48.7 0.6 5.0
Llama 3.1 8B 73.2 67.0 6.2 2.3 1.2 65.3 42.8 65.9 66.4 45.8 0.5 1.7
Qwen3 1.7B 49.0 53.4 9.8 2.3 0.3 62.9 42.4 54.3 61.9 46.6 0.7 2.0
Qwen3 4B 62.6 54.3 8.1 4.4 2.7 66.3 38.2 56.1 65.8 41.2 0.1 5.0
Qwen3 8B 52.2 57.4 4.1 2.1 0.4 51.4 54.5 58.7 52.5 57.5 0.4 1.6
Gemma 7B 64.4 42.3 8.5 5.0 2.2 54.0 46.3 45.1 52.6 46.0 0.8 3.4
Mistral 7B 68.6 65.2 6.8 2.5 1.2 63.9 44.7 65.2 65.6 47.9 0.7 2.2
Table 3: Comparison of TP and NLI bias scores for off-the-shelf (no debiasing) models. NLI bias scores are unstable and have limited correlation with TP bias scores. Bold scores represent cases where models appear to be unbiased on TP but display bias on NLI; italics scores represent cases where models appear to be unbiased on NLI but display bias on TP.
3.2 NLI Metric Selection
When comparing token probability and NLI bias scores, we notice that the overall behavior of NLI bias metrics is concerningly random. None of the metrics behave predictably with respect to token probability bias scores, and there is no obvious “best” metric. First, we observe that the threshold comparison and $\operatorname*{argmax}$ metrics $(M_{2},M_{3},M_{4},M_{10},M_{11})$ are not numerically comparable to TP and cannot be interpreted with as percentile scores where 50 is perfectly unbiased. The other six metrics seem to be at least roughly comparable to TP.
We want to understand if this poor correlation is caused by the relationship between TP and NLI bias scores, or if the NLI metrics we tested are generally brittle with high randomness. We acknowledge that token probability bias measurement has known issues and that the relationship between upstream bias metrics and downstream harms needs further study; however, we limit this analysis to comparison between TP and NLI as potential upstream metrics. For this analysis, we assume TP is an acceptable bias metric because it is a relatively established method in the literature, and we compare our proposed NLI metrics to the TP baseline. An ideal NLI metric should be at least somewhat linearly correlated with TP bias metrics. We expect that two metrics measuring this property on the same percentile scale would be correlated. If the correlation is weak or nonexistent, we conclude that one of the metrics is a poor measurement of the underlying property.
Metric $R^{2}$ Pearson $r$ $p$-value
$M_{1}$ 0.0681 0.261 0.1299
$M_{2}$ 0.0043 0.0655 0.7084
$M_{3}$ 0.0131 0.1146 0.512
$M_{4}$ 0.0434 0.2084 0.2295
$M_{5}$ 0.0001 0.0089 0.9594
$M_{6}$ 0.14 0.3741 0.0268
$M_{7}$ 0.0772 0.2778 0.1062
$M_{8}$ 0.0122 0.1103 0.5281
$M_{9}$ 0.1365 0.3695 0.0289
$M_{10}$ 0.0528 0.2297 0.1844
$M_{11}$ 0.0176 0.1327 0.4474
Table 4: $R^{2}$, Pearson $r$, and linear correlation $p$-values between TP bias scores and each NLI metric. Best correlation is bolded and second best is italicized.
The results of this correlation analysis are summarized in Table 4. All correlations are very weak, with a maximum $R^{2}$ value of $0.14$ for $M_{6}$ and a second-highest $R^{2}$ value of $0.1365$ for $M_{9}$. All other $R^{2}$ values are less than $0.1$. We also perform a two-sided Wald test with a $t$-distribution, with the null hypothesis that there is no relationship between TP and NLI metrics. The $p$-values for $M_{6}$ and $M_{9}$ are .0268.0268 and .0289.0289, respectively, indicating that the linear relationships between TP and $M_{6}$ and TP and $M_{9}$ are statistically significant at $\alpha=0.05$. All other $p$-values are greater than $0.1$, meaning the linear relationships are not statistically significant. These results confirm our intuition that $M_{6}$ and $M_{9}$ would be the best-performing metrics. However, both correlations are very weak, so we conclude that token probability and NLI, when formulated as percentile bias scores, do not seem to be measuring the same model behavior.
We also consider the linear correlations among different NLI bias scores. Using the same intuition as above, we argue that NLI bias metrics which are measuring the same thing should be somewhat correlated, and if 2 metrics have little to no correlation, one or both of them is noisy or brittle. The pairwise $R^{2}$ values for all 11 NLI metrics are shown in Fig. 1. Out of 55 pairwise combinations of NLI metrics, we find only ten with $R^{2}\geq 0.5$. Of these pairs, most unsurprisingly correspond to similarly formulated metrics. We also observe that the threshold comparison and $\operatorname*{argmax}$ metrics $(M_{2},M_{3},M_{4},M_{10},M_{11})$, while not well-correlated with TP, are moderately correlated with each other. The most interesting result here is the very strong correlation between $M_{6}$ and $M_{9}$, which were also the most closely correlated with TP. However, the overall behavior of the tested NLI metrics seems to be brittle and hard to predict. Given this finding, we conclude that $M_{6}$ and $M_{9}$ are the best NLI metrics, with very little statistical difference between them. NLI bias metrics seem to be generally noisy and we do not recommend their use as a replacement for TP bias metrics. However, when NLI bias metrics are necessary, we recommend that practitioners use either $M_{6}$ or $M_{9}$,, with the caveat that these metrics skew slightly lower than TP bias scores.
3.3 TP and NLI Results on Debiased Models
Next, we consider the debiased versions of BERT, RoBERTa, and GPT2 models. These results are summarized in Table 5. The concerning trend here is that NLI bias metrics often get higher after debiasing, while TP bias metrics uniformly move down. These instances are bolded in the table; the trend is particularly evident for BERT Base Cased, BERT Large Uncased, and RoBERTa Large. This suggests that NLI bias metrics respond less predictably than TP metrics to debiasing via continued pretraining.
Model TP R. TP N. TP T. $M_{6}$ R. $M_{6}$ N. $M_{6}$ T. $M_{9}$ R. $M_{9}$ N. $M_{9}$ T.
BERT B. Un. 74.49 45.71 41.05 54.66 41.57 44.92 57.10 43.81 47.05
BERT B. C. 64.40 61.67 57.81 40.71 41.68 45.37 42.87 46.22 47.23
BERT Lg. Un. 64.14 53.10 43.19 43.50 47.74 50.39 45.86 49.74 51.60
BERT Lg. C. 70.69 58.52 56.94 46.42 38.23 39.95 48.67 41.78 41.90
RoBERTa B. 69.18 64.33 54.34 52.27 44.94 44.11 59.05 52.46 51.62
RoBERTa Lg. 71.09 57.19 58.45 36.90 46.94 44.27 41.44 48.5 45.14
GPT2 68.27 49.82 45.11 47.83 36.66 35.09 53.00 40.25 41.26
GPT2 Med. 55.83 44.29 38.73 39.72 35.70 38.16 45.27 39.70 46.28
GPT2 XL 66.15 65.33 36.73 50.69 37.40 34.26 56.33 42.64 35.37